OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd? A.1/3 B.1/2 C.1/4 D.1/8

OpenStudy (anonymous):

How many odd numbers are between 1 and 10?

OpenStudy (anonymous):

answer is B. using what romero stated

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

both numbers are odd, not one number is odd the probability that kyle picks and odd number is \(\frac{1}{2}\) and the probability that lucy picks an odd number is also \(\frac{1}{2}\) so the probability that they both pick odd numbers is \(\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\)